OpenStudy (anonymous):

Solve for x: (x+20)/2 = 3x x = 10 x = 4 x = −14 x = −17

OpenStudy (anonymous):

i don't want to be rude, but shouldn't you try to answer them yourself? if you can't do it tell us and we'll help in the part that you are stuck :)

OpenStudy (ajprincess):

x+20=3x*2 x+20=6x 20=6x-x 20=5x x=20/5 x=4

OpenStudy (diyadiya):

Multiply both sides by 2 x+20=6x Subtract x from both sides 20=6x-x 20=5x divide by 5 x= ?

OpenStudy (diyadiya):

Someone already gave the answer ;/